# Local invariants on an ICIS

### v1

@@ Abstract
This work investigates invariants, including the GSV-index, local Euler obstruction, and Brasselet number, within the context of isolated complete intersection singularities (ICIS). The aim is to establish connections between these invariants, allowing for the extraction of topological and geometric information through algebraic methods. The Bruce-Roberts number plays a central role in achieving this objective, since is defined by algebraic tools.
Introduction
In the field of singularity theory, comprehending and quantifying the properties of singularities are crucial endeavors. One approach involves exploring invariants associated with these singularities. At the core of our inquiry lies the notion of Milnor number.
Let $(X,0)=(f^{-1}(0),0)$, where $f:(\mathbb{C}^{n+1},0)\rightarrow(\mathbb{C},0)$, be a germ of a hypersurface with an isolated singularity at the origin. Milnor proved in [18] that the Milnor fiber $X_{t}=f^{-1}(t)\cap\mathbb{B}_{\varepsilon}$ has the homotopy type of a bouquet of spheres with a real dimension $n=\text{dim}(X)$. The Milnor number is the count of these spheres and is equal to the complex dimension of the quotient $\frac{\mathcal{O}_{n}}{Jf}$, where $\mathcal{O}_{n}$ denotes the ring of analytic function-germs at the origin and $J(f)$ is the Jacobian ideal of $f$.
In Section 1, we introduce the concept of complete intersection singularities (ICIS). These singularities emerge as a natural extension of hypersurfaces, characterized by being loci of map germs that preserve the expected codimension. For such singularities, the Milnor number is also well-defined, as guaranteed by [12]. The Lê-Greul formula [14, 9] provides an algebraic expression for computing the Milnor number, concepts that we revisit.
@@ Introduction
Furthermore, certain generalizations aim to uncover insights when a function is defined over a singular germ complex space $(X,0)$. This is exemplified by the Bruce-Roberts number, introduced in [5] and revisited in Section 3. Defined through an algebraic approach, this number allows us to utilize the Singular software for computation. Our primary aim is to establish connections between the discussed invariants, providing a means to extract topological and geometric information by computing them via the Singular program.
@@ 1 The Milnor number of an ICIS
Complete intersections represent a natural extension of hypersurfaces in the following manner: a hypersurface singularity $(X,0)$ is a germ set in $\mathbb{C}^{n+1}$ defined as the locus where a single non-constant holomorphic function $f:(\mathbb{C}^{n+1},0)\rightarrow(\mathbb{C},0)$ vanishes. This condition implies that $(X,0)$ must possess dimension $n$. Now, if we consider $k$ holomorphic functions $f=(f_{1},\ldots,f_{k}):(\mathbb{C}^{n+k},0)\rightarrow(\mathbb{C}^{k},0)$, it is generally not the case that the locus $(X,0)$ of $f$ also has dimension $n$. However, when this does occur, we refer to $(X,0)$ as a complete intersection.
@@
A germ complex space $(X,0)\subset(\mathbb{C}^{n},0)$ is termed a complete intersection if the minimal number of generators of its ideal $\mathcal{I}(X,0)$ in $\mathcal{O}_{n}$ is $n-dim(X,0)$. An isolated complete intersection singularity (ICIS) refers to a complete intersection $(X,0)$ such that its singular locus, denoted as $Sing(X)$, coincides with the origin.
A comprehensive understanding of a singularity necessitates observing its transformation into less degenerate types under deformation. This involves examining a continuous family of varieties parametrized by a parameter space, where the original singularity emerges at a certain parameter value. Essentially, to comprehend the essence of a singularity, one must explore a spectrum of perspectives, typically represented by a 1-parameter family. A semiuniversal deformation, a special type of deformation, is characterized by its local simplicity. More precisely, it is a deformation that is universal among those exhibiting the same first-order behavior.
To facilitate our discussion, let $(X,0)\subset(\mathbb{C}^{n},0)$ denote an ICIS, with its defining ideal $\mathcal{I}(X,0)={\langle f_{1},\ldots,f_{k}\rangle}$, and $\{e_{i}\}_{i=1}^{n}$ representing the canonical basis of $\mathbb{C}^{n}$. The Tjurina number corresponds to the dimension of the base of a semiuniversal deformation of $(X,0)$ and on an ICIS is computed as:
@@
An algebraic formula for $\mu(X,0)$ has been independently provided by the authors Greul [9] and Lê [14], which is commonly referred to as the Lê-Greul formula. Let $f=(f_{1},\ldots,f_{k}):(\mathbb{C}^{n},0)\rightarrow(\mathbb{C}^{k},0)$ be a germ of holomorphic mapping defining an ICIS $(X,0)=(f^{-1}(0),0)$, and let $g=(f_{1},\ldots,f_{k-1})$. Denote $(Y,0)=(g^{-1}(0),0)$ and $J_{k}(f)$ as the ideal generated by all $k\times k$ minors of the Jacobian matrix of $f$. Then,
@@ 3 The Bruce-Roberts number
In this sense, there exists an invariant that maintains the aforementioned property of the Milnor number on germs of functions defined on singular germ complex spaces $(X,0)\subset(\mathbb{C}^{n},0)$, namely, its finiteness is equivalent to $f$ having an isolated singularity on the analytic space $(X,0)$, which in turn is equivalent to $f$ being $\mathcal{R}_{X}$-finitely determined, where $\mathcal{R}_{X}$ is the subgroup of $\mathcal{R}$ consisting of coordinate changes preserving $X$. This is the case of the number introduced by Bruce and Roberts in [5].
@@
Even though there is no explicit form for the generators of $\Theta_{X}$ when $(X,0)$ is any analytic variety, we can utilize software tools like Singular to compute these generators. Below is the step-by-step process, as formulated in [15].
@@
Based on these results, we can establish connections with other discussed invariants. The following proposition allows us to compute these invariants using the algebraic formulas provided by the Bruce-Roberts number and the relative Bruce-Roberts number.
@@
Proof. Consider a $1$-parameter deformation of $(X,0)$ given by a family $(\mathfrak{X},0)$ of smoothings, with the projection $\pi:(\mathfrak{X},0)\rightarrow(\mathbb{C},0)$.
@@
${\rm Eu}_{g,\mathfrak{X}\cap\pi^{-1}(\delta)}(x)=\mu(X,0)+\mu(X\cap f^{-1}(0),0),$
since $X=\mathfrak{X}\cap\pi^{-1}(0)$.
@@
$\overline{\mu_{BR}}(f,X)=\mu(f)+{\rm Eu}_{g,\mathfrak{X}\cap\pi^{-1}(\delta)}(x)-\tau(X,0).$
The desired expressions follow directly from the relations obtained in Theorem 2.7. ∎
@@
Combining the findings from [19], we additionally observe:
@@
    $\mu(X,0)+(-1)^{d-1}({\rm B}_{f,X}(0)-1)-\tau(X,0)$.

### v3

# Bruce-Roberts numbers and indices of vector fields on an ICIS

@@ Abstract
This work investigates invariants, including the GSV-index, the local Euler obstruction, and the Brasselet number, within the context of isolated complete intersection singularities (ICIS). The goal is to forge connections among these invariants, facilitating the extraction of both topological and geometric insights through algebraic methods. The Bruce-Roberts number assumes a central role in realizing this objective, given its definition via algebraic tools.
keywords:
local invariants, ICIS, Bruce-Roberts number, GSV-index, Euler obstruction
1 Introduction
In the field of singularity theory, comprehending and quantifying the properties of singularities are crucial endeavors. One approach involves exploring invariants associated with these singularities. Throughout this work, we elucidate the definitions and properties of local invariants, highlighting their role in characterizing singularities in complex analytic varieties.
At the core of our inquiry lies the notion of Milnor number. Let $(X,0)=(f^{-1}(0),0)$, where $f:(\mathbb{C}^{n+1},0)\rightarrow(\mathbb{C},0)$, be the germ of a hypersurface with an isolated singularity at the origin. Milnor proved in [27] that the fiber $X_{t}=f^{-1}(t)\cap\mathbb{B}_{\varepsilon}$, for a sufficiently small ball $\mathbb{B}_{\varepsilon}$ centered at 00, has the homotopy type of a bouquet of spheres with a real dimension $n=\text{dim}(X)$. The Milnor number is the count of these spheres and is equal to the length of the analytic algebra $\frac{\mathcal{O}_{n}}{Jf}$, where $\mathcal{O}_{n}$ denotes the ring of analytic function-germs at the origin and $J(f)$ is the Jacobian ideal of $f$.
In Section 2, we recall the concept of complete intersection singularities (ICIS). These singularities emerge as a natural extension of hypersurfaces, characterized by being loci of map germs that preserve the expected codimension. For such singularities, the Milnor number is defined for Hamm in [19]. The Lê-Greul formula [21, 14] provides an algebraic expression for computing the Milnor number, concepts that we revisit.
@@ 1 Introduction
Furthermore, certain generalizations aim to uncover insights when a function is defined over a singular complex space germ $(X,0)$. This is exemplified by the Bruce-Roberts number, introduced in [8] and revisited in Section 4. Defined through an algebraic approach, this number allows us to utilize the Singular software [10] for computation. Our primary aim is to establish connections between the discussed invariants, providing a means to extract topological and geometric information by computing them via the Singular program.
@@ 2 The Milnor number of an ICIS
Complete intersections represent a natural extension of hypersurfaces in the following manner: a hypersurface singularity $(X,0)$ is a germ set in $\mathbb{C}^{n+1}$ defined as the locus where a single non-constant holomorphic function $f:(\mathbb{C}^{n+1},0)\rightarrow(\mathbb{C},0)$ vanishes. This condition implies that $(X,0)$ must possess dimension $n$. Now, if we consider $k$ holomorphic functions $f=(f_{1},\ldots,f_{k}):(\mathbb{C}^{n+k},0)\rightarrow(\mathbb{C}^{k},0)$, it is generally not the case that the zero locus $(X,0)$ of $f$ also has dimension $n$. However, when this does occur, we refer to $(X,0)$ as a complete intersection.
@@
A complex space germ $(X,0)\subset(\mathbb{C}^{n},0)$ is termed a complete intersection if the minimal number of generators of its ideal $\mathcal{I}(X,0)$ in $\mathcal{O}_{n}$ is $n-dim(X,0)$. An isolated complete intersection singularity (ICIS) refers to a complete intersection $(X,0)$ such that its singular locus, denoted as $Sing(X)$, coincides with the origin.
A comprehensive understanding of a singularity involves observing its transformation into less degenerate types under deformation. This requires examining a continuous family of varieties parametrized, where the original singularity emerges at a certain parameter value. Essentially, to comprehend the essence of a singularity, one must explore a spectrum of perspectives, typically represented by a 1-parameter family. A semiuniversal deformation, a special type of deformation, is characterized by its local simplicity. Roughly speaking, it is a deformation that is universal among those exhibiting the same first-order behavior.
To facilitate our discussion, let $(X,0)\subset(\mathbb{C}^{n},0)$ denote an ICIS, with its defining ideal $\mathcal{I}(X,0)={\langle f_{1},\ldots,f_{k}\rangle}$, and $\{e_{i}\}_{i=1}^{n}$ representing the canonical basis of $\mathbb{C}^{n}$. The Tjurina number corresponds to the dimension of the base of a semiuniversal deformation of $(X,0)$ and on an ICIS can be computed as follows. This equality is established as Theorem 1.6 in [15].
@@
An algebraic formula for $\mu(X,0)$ has been independently provided by Greul [14] and Lê [21], which is commonly referred to as the Lê-Greul formula. Let $f=(f_{1},\ldots,f_{k}):(\mathbb{C}^{n},0)\rightarrow(\mathbb{C}^{k},0)$ be a germ of holomorphic mapping defining an ICIS $(X,0)=(f^{-1}(0),0)$, and let $g=(f_{1},\ldots,f_{k-1})$. Denote $(Y,0)=(g^{-1}(0),0)$ and $J_{k}(f)$ as the ideal generated by all $k\times k$ minors of the Jacobian matrix of $f$. Then,
@@ 3.1 The GSV-index
Another noteworthy property of the Poincaré–Hopf index on a germ of a holomorphic vector field $v=\sum_{i=1}^{n}h_{i}\frac{\partial}{\partial z_{i}}$ on $\mathbb{C}^{n+1}$ with an isolated singularity at 00 is that the local index equals the integer:
$\dim_{\mathbb{C}}\dfrac{\mathcal{O}_{n}}{\left<h_{1},\dots,h_{n}\right>},$
where $h_{1},\dots,h_{n}\right>$ is the ideal generated by the components of $v$ in the ring $\mathcal{O}_{n}$. When considering $v$ as the complex conjugate of the gradient vector field of a function $f:(\mathbb{C}^{n+1},0)\rightarrow(\mathbb{C},0)$, it follows that the Poincaré-Hopf index of $v$ aligns with the Milnor number associated with the hypersurface defined by $f$. Hence, the GSV-index can be interpreted as an extension or generalization of the concept of the Milnor number.
Moreover, this characterization motivates the search for algebraic formulas for the index of vector fields on singular varieties. The homological index, defined by Gómez-Mont in [13], provides a solution to this inquiry. It considers an isolated singularity germ $(X,0)$ of arbitrary dimension, accompanied by a holomorphic vector field on $X$ that is singular only at 00. Within this context, we employ Kähler differentials on $X$ and construct a Koszul complex $(\Omega^{\bullet}_{X,0},v)$:
$0\rightarrow\Omega^{n}_{X,0}\rightarrow\Omega^{n-1}_{X,0}\rightarrow\cdots\rightarrow\mathcal{O}_{X,0}\rightarrow 0,$
where the arrows denote contraction of forms by the vector field $v$. The homological index of $v$ is then defined as the Euler characteristic of this complex.
Definition 3.3.
The homological index, denoted as $ind_{hom}(v,X,0)$, of the holomorphic vector field $v$ on $(X,0)$ is defined as:
$ind_{hom}(v,X,0)=\sum_{i=0}^{n}(-1)^{i}h_{i}(\Omega^{\bullet}_{X,0},v),$
where $h_{i}(\Omega^{\bullet}_{X,0},v)$ represents the dimension of the $i$-th homology group of the complex as a vector space over $\mathbb{C}$.
The authors in [2] have demonstrated that the homological index corresponds to the GSV-index for an ICIS. With this understanding, we refer to the GSV-index in our discussion. It is worth noting that while the homological index is defined for vector fields on arbitrary isolated normal singularity germs, the GSV-index is specifically defined on complete intersection germs.
@@ 3.2 The local Euler obstruction
The determination of the local Euler obstruction often relies on topological methods, including the application of the Euler characteristic. Brasselet, Lê, and Seade have outlined a formula for this purpose in their work [4].
@@
Remark 3.9.
As previously mentioned, Theorem 3.8 was originally formulated for isolated determinantal singularities. In this case, the so-called Poincaré-Hopf-Nash index is well-defined, which, in the case of ICIS, coincides with the GSV-index.
@@ 4 The Bruce-Roberts number
In this sense, there exists an invariant that maintains the aforementioned property of the Milnor number on germs of functions defined on singular complex space germs $(X,0)\subset(\mathbb{C}^{n},0)$, namely, its finiteness is equivalent to $f$ having an isolated singularity on the analytic space $(X,0)$, which in turn is equivalent to $f$ being $\mathcal{R}_{X}$-finitely determined, where $\mathcal{R}_{X}$ is the subgroup of $\mathcal{R}$ consisting of coordinate changes preserving $X$. This is the case of the number introduced by Bruce and Roberts in [8].
@@
Even though there is no explicit form for the generators of $\Theta_{X}$ when $(X,0)$ is an analytic variety, we can utilize software tools like Singular to compute these generators. Below is the step-by-step process, as formulated in [22].
@@
Based on these results, we can establish connections with other discussed invariants. In the literature, connections between the Bruce-Roberts number and the local Euler obstruction are documented, as in [16], and with the Chern Obstruction as demonstrated in [24]. The following proposition enables us to compute these invariants using the algebraic formulas provided by the Bruce-Roberts number and the relative Bruce-Roberts number.
@@
Proof 4.5.
Consider a $1$-parameter deformation of $(X,0)$ given by a family $(\mathfrak{X},0)$ of smoothings, with the projection $\pi:(\mathfrak{X},0)\rightarrow(\mathbb{C},0)$.
@@
${\rm Eu}_{g,\mathfrak{X}\cap\pi^{-1}(\delta)}(x)=\mu(X,0)+\mu(X\cap f^{-1}(0),0).$
@@
$\overline{\mu_{BR}}(f,X)={\rm Eu}_{g,\mathfrak{X}\cap\pi^{-1}(\delta)}(x)-\tau(X,0).$
The desired expressions follow directly from the relations obtained in Theorem 3.8.
@@
In the context of hypersurfaces, the authors of [28] and [25] independently presented relationships among these Milnor numbers. Combining the findings, we additionally observe:
@@
    $\mu(f)+\mu(X,0)+\mu(X\cap p^{-1}(0),0)+(-1)^{d}{\rm Eu}_{f,X}(0)-\tau(X,0)$, where $p:\mathbb{C}^{n}\rightarrow\mathbb{C}$ is a linear generic function with respect to $X$;
@@
    $\mu(f)+\mu(X,0)+(-1)^{d-1}({\rm B}_{f,X}(0)-1)-\tau(X,0)$.
In their work [1], the authors delve into the investigation of the so-called Bruce-Roberts Tjurina number of the function $f$ concerning the variety $(X,0)$, particularly when there exists a coordinate system such that both $f$ and $X$ are quasihomogeneous with respect to the same positive rational weights. The Bruce-Roberts Tjurina number is formally defined as follows:
$\tau_{BR}(f,X):=\dim_{\mathbb{C}}\dfrac{\mathcal{O}n}{df(\Theta_{X})+\left<f\right>}.$
This quantity characterizes the infinitesimal deformations of $f$ under $\mathcal{K}_{X}$-equivalence, representing diffeomorphisms that preserve $X$ and allow for the multiplication of $f$ by units in $\mathcal{O}_{n}$. The authors have shown that the equality $\mu_{BR}(f,X)=\tau_{BR}(f,X)$ is indicative of the relative quasihomogeneity of the pair $(f,X)$. In such cases, the Bruce-Roberts Tjurina number can also be computed using the expressions provided above.
Acknowledgments